# Galaxy mass profiles with convolutional neural networks

### v1

@@ Abstract
Context. Determining the dynamical mass profiles of dispersion-supported galaxies is particularly challenging due to projection effects and the unknown shape of their velocity anisotropy profile.
Aims. Our goal is to develop a machine learning algorithm capable of recovering dynamical mass profiles of dispersion-supported galaxies from line-of-sight stellar data. Traditionally, this task relies on time-consuming methods that require profile parameterization and assume dynamical equilibrium and spherical symmetry.
Methods. We train a convolutional neural network model using various sets of cosmological hydrodynamical simulations of galaxies. By extracting projected stellar data from the simulated galaxies and feeding it into the model, we obtain the posterior distribution of the dynamical mass profile at ten different radii. Additionally, we evaluate the performance of existing literature mass estimators on our dataset.
Results. Our model achieves more accurate results than any literature mass estimator while also providing enclosed mass estimates at radii where no previous estimators exist. We confirm that the posterior distributions produced by the model are well-calibrated, ensuring they provide meaningful uncertainties. However, issues remain, as the method loses performance when trained on one set of simulations and applied to another, highlighting the importance of improving the generalization of ML methods trained on specific galaxy simulations.
@@
Understanding the dynamical mass distribution within galaxies is crucial for testing the cold dark matter paradigm. In rotation-supported galaxies, mass is inferred from rotational velocities, whereas dispersion-supported galaxies require kinematic studies based on spectroscopic data to trace stellar motions. However, these measurements are subject to projection effects, making it difficult to reconstruct the full three-dimensional structure of the galaxy.
@@
Dynamical modelling provides a robust approach to address these challenges. Advanced techniques, often combined with Monte Carlo simulations, yield robust mass estimates with well-calibrated uncertainties. However, these models are still affected by the mass-anisotropy degeneracy; unknown velocity anisotropies can bias the inferred mass distribution, with some methods being more sensitive to these effects than others. Moreover, the computational expense and the need for careful system modelling further complicate the process.
@@
In contrast with dynamical modelling, several authors (e.g., Walker et al., 2009; Wolf et al., 2010; Amorisco & Evans, 2012; Campbell et al., 2017; Errani et al., 2018) have developed simple formulae to estimate the dynamical mass of galaxies enclosed within specific radii, where they have found velocity anisotropy and/or other factors to minimise uncertainty. These estimators rely on the line-of-sight velocity dispersion of the stellar component and on the half-light radius, aiming to obtain an unbiased mass estimate that holds for a broad range of galaxy properties and that is minimally affected by line-of-sight projection effects.
@@
In recent years, machine learning has proven to be an effective tool for analysing complex data and uncovering underlying patterns. These methods have been applied to various astrophysical problems, including dynamical mass estimation and density profile determination using line-of-sight data. For example, de los Rios et al. (2023) used convolutional neural networks (CNNs) to analyse mock photometry and interferometry images from cosmological hydrodynamical simulations of spiral galaxies, successfully inferring 20 points of their dynamical mass profiles. Similarly, Expósito-Márquez et al. (2023) employed 2-d probability distribution functions of projected stellar positions and kinematics as inputs to a CNN to estimate the inner slopes (150 pc from the centre) of dark matter density profiles in dwarf galaxies. This approach resulted in a model which was able to differentiate between cusps and cores in cosmological hydrodynamical simulations. Ho et al. (2019) applied a CNN model with a similar input structure based on probability distribution functions of positions and velocities of galaxies to estimate the dynamical masses of galaxy clusters. Also, Nguyen et al. (2023) developed a graph neural network capable of accurately recovering the dark matter density profiles of mock spherical, dynamically equilibrated dwarf galaxies.
@@
In this paper we build a CNN based on that used in Expósito-Márquez et al. (2023), and we train it on a mixed dataset of cosmological hydrodynamical simulations of dispersion-supported galaxies spanning a broad range of stellar mass and other galactic properties. Our model uses projected stellar information to estimate the dynamical mass profile of galaxies providing the posterior distribution of the values of ten points of the dynamical mass profile of the galaxy. We compare the capabilities of our model with five existing literature mass estimators (Walker et al., 2009; Wolf et al., 2010; Amorisco & Evans, 2012; Campbell et al., 2017; Errani et al., 2018). We find our model provides improved scatter in the estimation of dynamical mass, on top of its ability to provide new points of the mass profile at radii where no literature mass estimators have been proposed.
@@
The structure of this article is the following: In Section 2 we describe the simulations used to test the literature mass estimators and train the CNN, and also the process we follow to prepare the dataset to be inputted into the model. In Section 3 we introduce the existing mass estimators in the literature, shortly explaining their derivation and assumptions. We describe our CNN model in Section 4, where we show its architecture and explain the structure of inputs and outputs. In Section 5 we present the results of our work. First in Section 5.1 we show the performance of literature mass estimators on our dataset of simulated galaxies and then in Section 5.2 we compare with the results of our model. We summarise our conclusions in Section 6.
@@ 2 Simulation Dataset
Training a CNN to estimate galaxy dynamical mass requires an extensive dataset of realistic simulated galaxies from which we can extract stellar data to use as input for the network, as well as calculate the enclosed mass profile we aim to estimate. We decide to use cosmological hydrodynamical simulations over idealised simulations, choosing to train our model on a more realistic dataset for which we have limited sampling capability, rather than on a less realistic dataset that we can generate and sample faster and more uniformly.
Since we are restricted to a dataset of previously run simulations, we aim for it to sample the space of galaxy properties as completely as possible, providing a large enough number of galaxies for the study. We find that a set of simulations based on the NIHAO project (Wang et al., 2015) fulfils these requirements. In order to assess possible biases in the model we also select other suites of simulations to compare the performance of the model. For that, we also make use of publicly available data for the AURIGA (Grand et al., 2024) and the FIRE-2 (Wetzel et al., 2023) simulations.
@@ 2.1 The NIHAO project
The NIHAO project consists in a series of cosmological hydrodynamical zoom-in simulations run with the cosmology parameters of Planck Collaboration et al. (2014): H = 100h km s Mpc with h = 0.671, = 0.3175, = 0.6824, = 0.049 and = 0.8344.
The galaxy formation model includes ultraviolet heating, ionisation and metal cooling (Shen et al., 2010). Star formation and feedback follow the model used in the Making Galaxies In a Cosmological Context simulations (MaGICC, Stinson et al. 2013), reproducing galaxy scaling relations over a wide mass range (Brook et al., 2012). The density threshold for star formation is and a Chabrier (2003) initial mass function is used. Energy from stars is injected into the ISM, modelled using blast-wave supernova feedback (Stinson et al., 2006) and early stellar feedback from massive stars. Particle masses and force softenings allow the mass profile to be resolved to below 1 of the virial radius, which ensures that galaxy half-light radii are well resolved with spatial resolution varying from 100 pc through 800 pc for the most massive galaxies.
@@ 2.2 The AURIGA project
The galaxy formation model implemented in simulations from the AURIGA project includes magnetic fields, primordial line metal cooling which takes account self-shielding, stellar feedback and thermal feedback from radio and quasar accretion modes of black holes. Star formation is regulated by the Kennicut-Schmidt law (Schmidt, 1959) and a Chabrier (2003) initial mass function. The resolution of publicly available simulations ranges from 200 pc to 400 pc.
@@ 2.3 The FIRE project
The FIRE project includes simulations with different cosmologies. Some of them follow Planck Collaboration et al. (2014), and others use the cosmology from the AGORA project (Kim et al., 2014): H = 100h km s Mpc with h = 0.702, = 0.272, = 0.728, = 0.0455 and = 0.807. Other simulations adopt the cosmology from the AGORA project.
The FIRE suite galaxy formation model implements stellar and super-massive black hole feedback, magnetic fields, cosmic rays, photo-ionization and photo-heating. They reach very high spatial resolution with softening lengths in the order of 10 parsecs.
The FIRE simulations, particularly the FIRE-2 suite, have been shown to effectively reproduce several key properties of real galaxies. Hopkins et al. (2018a) demonstrated that FIRE-2 galaxies align with observed morphologies, interstellar medium structure, star formation histories, rotation curves, metallicities and stellar masses.
@@ 2.4 Training Dataset
Our dataset comprises 496 zoom-in cosmological simulations extracted from the NIHAO project, 66 zoom-in simulations from the AURIGA project and 17 zoom-in simulations from the FIRE project. Besides the fiducial NIHAO dataset, we include simulations of NIHAO galaxies with different stellar formation density thresholds (Dutton et al., 2020), NIHAO galaxies including black holes (Blank et al., 2019) and high resolution runs of 6 NIHAO simulations (Buck et al., 2019). Due to the limited number of FIRE simulations, galaxies from this suite will be used only for testing the model and not for training.
@@
For all simulations in the NIHAO+AURIGA dataset, we select all galaxies with stellar masses ranging from to that contain at least 50 star particles and whose mass fraction in high resolution particles is higher than 95%. Out of such galaxies, we select the ones that are considered dispersion-supported systems according to the criterion . , introduced in Correa et al. (2017), measures the kinetic energy fraction invested in ordered rotation, and is defined as
@@
where is the total kinetic energy of stars, and the sum is over all star particles within 30 kpc of the centre of the galaxy and with positive values of the specific angular momentum () along the direction of the total angular momentum of the stellar component of the galaxy. is the mass of each stellar particle and is the distance from the star to the rotation axis of the stellar component. We further improve the quality of the dataset by manually removing ongoing mergers and heavily disrupted galaxies identified through visual inspection. Applying these filters results in a sample of 1976 galaxies from the NIHAO project, 1066 from the AURIGA project, and 192 from the FIRE project.
@@
We sample stars randomly from each galaxy with a sample size between 200 and , depending on the total number of star particles in the original galaxy. Then we project the stellar data of galaxies along 64 lines-of-sight uniformly distributed along the unit sphere. We carry on this process for galaxies of all three simulations in order to construct datasets where we can test the performance of models over all lines-of-sight. For each projection we also calculate the projected half-light radius (half-stellar number111To avoid relying on any stellar flux model, we assume that all stars have the same luminosity. Consequently, when we mention the half-light radius of the simulated galaxies, we are referring to their half-stellar number radius.) and the line-of-sight velocity dispersion. Since the datasets from NIHAO and AURIGA simulations are fairly large (105), we decide to cut them by selecting the same number of projections from each suite for each bin of dynamical mass enclosed within the half-light radius, dropping the excess projections. We use 30 bins between dynamical enclosed masses 107.5 and 1011. This process allows for a well balanced training of the neural network on a mixed dataset of AURIGA and NIHAO galaxies. We show the resulting mass distributions in Fig. 1.
In Fig. 2 we show the contours of the distribution in the space of stellar mass - projected half-light radius for all simulations.
@@
The final training dataset consists in a total of 32412 projections of galaxies from the NIHAO project and the same number from the AURIGA project. We also obtain a dataset of 10398 projections of galaxies from the FIRE project within the same range of dynamical masses (within the projected half-light radius) as the NIHAO and AURIGA datasets. When training our neural network we make sure that all the projections of an individual galaxy lay either within the training set or the validation set, avoiding the validation of the model on data extracted from a galaxy that has also been used to train the model.
@@ 3 Mass Estimators
Several works in the literature have developed simple formulae for calculating galaxy dynamical masses enclosed within different radii from projected quantities only. These estimators make use of the projected half-light radius of the galaxy and the line-of-sight velocity dispersion of the stellar component either in all the galaxy, at a specific distance from the centre or within a given radius. We introduce here the most common mass estimators that we test on our simulations in Sec. 5.1.
The first of such mass estimators is that of Peñarrubia et al. (2008). They find that in the case of isotropic () stellar components following a King profile inside a Navarro et al. (1997) halo, the error of using the central stellar component’s projected velocity dispersion for calculating the dynamical mass of galaxies is minimised within the King core radius (). Their mass estimator takes the form .
Walker et al. (2009) make the less restrictive assumption that the anisotropy profile is flat throughout all the galaxy, and that the stellar component follows a Plummer spherical profile. They argue that the error on mass estimation introduced by the constant anisotropy is small when using the estimator that uses the line-of-sight stellar velocity dispersion over the whole galaxy.
@@
For stellar components modelled using the Michie-King distribution function, Amorisco & Evans (2012) find .
Campbell et al. (2017) used the APOSTLE simulations (Sawala et al., 2016) to derive a mass estimator for larger radii .
For a similar radius, Errani et al. (2018) found that minimises the errors introduced by the shape of the tidally-stripped dark matter halo (cusp or core) in which lays a Plummer sphere stellar component.
Some mass estimators have been tested on cosmological simulations. Apart from their own estimator, Campbell et al. (2017) tested Walker’s and Wolf’s mass estimators on fully cosmological hydrodynamical simulations from the APOSTLE project. González-Samaniego et al. (2017) tested Walker’s, Wolf’s and Campbell’s mass estimators on 10 dwarf galaxy cosmological hydrodynamical simulations from the FIRE-2 project (Hopkins et al., 2018b). Both studies obtained different results for the biases and specially for the scatters of the three mass estimators.

### v4

# Dispersion-supported galaxy mass profiles with convolutional neural networks

@@ Abstract
Context.
Aims. Determining the dynamical mass profiles of dispersion-supported galaxies is particularly challenging due to projection effects and the unknown shape of their velocity anisotropy profile. Traditionally, this task relies on time-consuming methods that require profile parameterisation and the assumption of dynamical equilibrium and spherical symmetry.
Methods. Our goal is to develop a machine-learning algorithm capable of recovering dynamical mass profiles of dispersion-supported galaxies from line-of-sight stellar data.
Results. We trained a convolutional neural network model using various sets of cosmological hydrodynamical simulations of galaxies. By extracting projected stellar data from the simulated galaxies and feeding them into the model, we obtained the posterior distribution of the dynamical mass profile at ten different radii. Additionally, we evaluated the performance of existing literature mass estimators on our dataset.
Conclusions. Our model achieves more accurate results than any literature mass estimator while also providing enclosed mass estimates at radii where no previous estimators exist. We confirm that the posterior distributions produced by the model are well calibrated, ensuring they provide meaningful uncertainties. However, issues remain: the method’s performance is less good when trained on one set of simulations and applied to another, highlighting the importance of improving the generalisation of machine-learning methods trained on specific galaxy simulations.
@@
Understanding the dynamical mass distribution within galaxies is crucial for testing the cold dark matter paradigm. In rotation-supported galaxies, mass is inferred from rotational velocities, whereas dispersion-supported galaxies require kinematic studies based on spectroscopic data to trace stellar motions. These measurements are subject to projection effects, making it difficult to reconstruct the full 3D structure of the galaxy.
@@
Dynamical modelling provides a robust approach to addressing these challenges. Advanced techniques, often combined with Monte Carlo simulations, yield robust mass estimates with well-calibrated uncertainties. However, these models are still affected by the mass-anisotropy degeneracy; unknown velocity anisotropies can bias the inferred mass distribution, with some methods more sensitive to these effects than others. Moreover, the computational expense and the need for careful system modelling further complicate the process.
@@
In contrast with dynamical modelling, simple formulae have been developed (e.g. Walker et al., 2009; Wolf et al., 2010; Amorisco & Evans, 2012; Campbell et al., 2017; Errani et al., 2018) to estimate the dynamical mass of galaxies enclosed within specific radii; in these approaches, velocity anisotropy and/or other factors have been found to minimise uncertainty. These estimators rely on the line-of-sight velocity dispersion of the stellar component and on the half-light radius, aiming to obtain an unbiased mass estimate that holds for a broad range of galaxy properties and is minimally affected by line-of-sight projection effects.
@@
In recent years, machine learning has proven to be an effective tool for analysing complex data and uncovering underlying patterns. These methods have been applied to various astrophysical problems, including dynamical mass estimation and density profile determination using line-of-sight data. For example, de los Rios et al. (2023) used convolutional neural networks (CNNs) to analyse mock photometry and interferometry images from cosmological hydrodynamical simulations of spiral galaxies, successfully inferring 20 points of their dynamical mass profiles. Similarly, Expósito-Márquez et al. (2023) employed 2D probability distribution functions (PDFs) of projected stellar positions and kinematics as inputs to a CNN to estimate the inner slopes (150 pc from the centre) of dark matter density profiles in dwarf galaxies. This approach resulted in a model that was able to differentiate between cusps and cores in cosmological hydrodynamical simulations. Ho et al. (2019) applied a CNN model with a similar input structure based on PDFs of the positions and velocities of galaxies to estimate the dynamical masses of galaxy clusters. Also, Nguyen et al. (2023) developed a graph neural network capable of accurately recovering the dark matter density profiles of mock spherical, dynamically equilibrated dwarf galaxies.
@@
In this study we built a CNN based on that used in Expósito-Márquez et al. (2023) and trained it on a mixed dataset of cosmological hydrodynamical simulations of dispersion-supported galaxies spanning a broad range of stellar masses and other galactic properties. Our model uses projected stellar information to estimate the dynamical mass profile of galaxies, providing the posterior distribution of the values of ten points of the dynamical mass profile of the galaxy. We compared the capabilities of our model with five existing literature mass estimators (Walker et al., 2009; Wolf et al., 2010; Amorisco & Evans, 2012; Campbell et al., 2017; Errani et al., 2018). We find our model provides improved scatter in the estimation of dynamical mass, on top of its ability to provide new points of the mass profile at radii where no literature mass estimators have been proposed.
@@
The structure of this article is the following: In Sect. 2 we describe the simulations used to test the literature mass estimators and train the CNN, and also the process we followed to prepare the dataset to be inputted into the model. In Sect. 3 we introduce the existing mass estimators in the literature, and briefly explaining their derivation and assumptions. We describe our CNN model in Sect. 4, where we also present its architecture and explain the structure of inputs and outputs. In Sect. 5 we present the results of our work. In Sect. 5.1 we show the performance of literature mass estimators on our dataset of simulated galaxies, and then in Sect. 5.2 we compare them with the results of our model. We summarise our conclusions in Sect. 6.
@@ 2 Simulation dataset
Training a CNN to estimate galaxy dynamical mass requires an extensive dataset of realistic simulated galaxies from which we can extract stellar data to use as input for the network, as well as calculate the enclosed mass profile we aim to estimate. We decided to use cosmological hydrodynamical simulations over idealised simulations, choosing to train our model on a more realistic dataset for which we have limited sampling capability, rather than on a less realistic dataset that we can generate and sample faster and more uniformly.
Since we are restricted to a dataset of previously run simulations, we aimed for it to sample the space of galaxy properties as completely as possible, providing a large enough number of galaxies for the study. We found a set of simulations based on the Numerical Investigation of a Hundred Astrophysical Objects (NIHAO; Wang et al., 2015) project that fulfils these requirements. In order to assess possible biases in the model, we also selected other suites of simulations to compare the performance of the model. For that, we also made use of publicly available data for the AURIGA (Grand et al., 2024) and the Feedback In Realistic Environments-2 (FIRE-2; Wetzel et al., 2023) simulations.
@@ 2.1 The NIHAO project
The NIHAO project consists in a series of cosmological hydrodynamical zoom-in simulations run with the cosmology parameters of Planck Collaboration et al. (2014): H = 100h km s Mpc with h = 0.671, = 0.3175, = 0.6824, = 0.049, and = 0.8344. The galaxy formation model includes ultraviolet heating, ionisation and metal cooling (Shen et al., 2010). Star formation and feedback follow the model used in the Making Galaxies In a Cosmological Context (MaGICC; Stinson et al., 2013) simulations, reproducing galaxy scaling relations over a wide mass range (Brook et al., 2012). The density threshold for star formation is and a Chabrier (2003) initial mass function is used. Energy from stars is injected into the interstellar medium, modelled using blast-wave supernova feedback (Stinson et al., 2006) and early stellar feedback from massive stars. Particle masses and force softenings allow the mass profile to be resolved to below 1 of the virial radius, which ensures that galaxy half-light radii are well resolved with spatial resolution varying from 100 pc through 800 pc for the most massive galaxies.
@@ 2.2 The AURIGA project
The galaxy formation model implemented in simulations from the AURIGA project includes magnetic fields, primordial line metal cooling that takes self-shielding into account, stellar feedback and thermal feedback from radio and quasar accretion modes of black holes. Star formation is regulated by the Kennicutt-Schmidt law (Schmidt, 1959) and a Chabrier (2003) initial mass function. The resolution of publicly available simulations ranges from 200 pc to 400 pc.
@@ 2.3 The FIRE project
The FIRE project includes simulations with different cosmologies. Some of them follow Planck Collaboration et al. (2014), and others adopt the cosmology from the Assembling Galaxies Of Resolved Anatomy (AGORA; Kim et al., 2014) project: H = 100h km s Mpc with h = 0.702, = 0.272, = 0.728, = 0.0455, and = 0.807.
The FIRE suite galaxy formation model implements stellar and super-massive black hole feedback, magnetic fields, cosmic rays, photo-ionisation and photo-heating. They reach very high spatial resolution with softening lengths in the order of 10 parsecs.
The FIRE simulations, particularly the FIRE-2 suite, have been shown to effectively reproduce several key properties of real galaxies. Hopkins et al. (2018a) demonstrate that FIRE-2 galaxies align with observed morphologies, interstellar medium structure, star formation histories, rotation curves, metallicities and stellar masses.
@@ 2.4 Training dataset
Our dataset comprises 496 zoom-in cosmological simulations extracted from the NIHAO project, 66 zoom-in simulations from the AURIGA project and 17 zoom-in simulations from the FIRE project. Besides the fiducial NIHAO dataset, we included simulations of NIHAO galaxies with different stellar formation density thresholds, (Dutton et al., 2020), NIHAO galaxies that have black holes (Blank et al., 2019), and high resolution runs of six NIHAO simulations (Buck et al., 2019). Due to the limited number of FIRE simulations, galaxies from this suite were only used to test the model and not for the training.
@@
For all simulations in the NIHAO+AURIGA dataset, we selected all galaxies with stellar masses ranging from to that contain at least 50 star particles and whose mass fraction in high resolution particles is larger than 95%. Out of these galaxies, we selected those that are considered dispersion-supported systems according to the criterion . The quantity was introduced in Correa et al. (2017). It measures the kinetic energy fraction invested in ordered rotation and is defined as
@@
where is the total kinetic energy of stars, and the sum is over all star particles within 30 kpc of the centre of the galaxy and with positive values of the specific angular momentum () along the direction of the total angular momentum of the stellar component of the galaxy. is the mass of each stellar particle, and is the distance from the star to the rotation axis of the stellar component. We further improved the quality of the dataset by manually removing ongoing mergers and heavily disrupted galaxies identified through visual inspection. Applying these filters resulted in a sample of 1976 galaxies from the NIHAO project, 1066 from the AURIGA project, and 260 from the FIRE project.
@@
We sampled stars randomly from each galaxy with a sample size between 200 and , depending on the total number of star particles in the original galaxy. For galaxies with fewer than 200 star particles, we selected all of them. Then we project the stellar data of galaxies along 64 lines of sight uniformly distributed along the unit sphere. We carried on this process for galaxies of all three simulations in order to construct datasets we could use to test the performance of models over all lines of sight. For each projection we also calculated the projected half-light radius (half-stellar number111To avoid relying on any stellar flux model, we assumed that all stars have the same luminosity. Consequently, when we mention the half-light radius of the simulated galaxies, we are referring to their half-stellar number radius.) and the line-of-sight velocity dispersion. Since the datasets from NIHAO and AURIGA simulations are fairly large (105), we decided to cut them by selecting the same number of projections from each suite for each bin of dynamical mass enclosed within the half-light radius, dropping the excess projections. We used 30 bins between dynamical enclosed masses 107.5 and 1011. This process enabled a well-balanced training of the neural network on a mixed dataset of AURIGA and NIHAO galaxies. We show the resulting mass distributions in Fig. 1. In Fig. 2 we show the contours of the distribution in the space of stellar mass–projected half-light radius for all simulations.
@@
The final training dataset consists in a total of 32344 projections of galaxies from the NIHAO project and the same number from the AURIGA project. We also obtained a dataset of 16640 projections of galaxies from the FIRE project within the same range of dynamical masses (within the projected half-light radius) as the NIHAO and AURIGA datasets. When training our neural network we made sure that all the projections of an individual galaxy lie within either the training set or the validation set, avoiding the validation of the model on data extracted from a galaxy that had also been used to train the model.
@@ 3 Mass estimators
Several works in the literature have developed simple formulae for calculating galaxy dynamical masses enclosed within different radii from projected quantities only. These estimators make use of the projected half-light radius () of the galaxy and the line-of-sight velocity dispersion of the stellar component (i) in all the galaxy, (ii) at a specific distance from the centre, or (iii) within a given radius. We introduce here the most common mass estimators that we tested on our simulations in Sect. 5.1.
The first such mass estimator is that of Peñarrubia et al. (2008). They find that in the case of isotropic () stellar components following a King profile inside a Navarro et al. (1997) halo, the error of using the central stellar component’s projected velocity dispersion for calculating the dynamical mass of galaxies is minimised within the King core radius (). Their mass estimator takes the form .
Walker et al. (2009) made the less restrictive assumption that the anisotropy profile is flat throughout all the galaxy, and that the stellar component follows a Plummer spherical profile. They argued that the error on mass estimation introduced by the constant anisotropy is small when using the estimator that uses the line-of-sight stellar velocity dispersion over the whole galaxy.
@@
For stellar components modelled using the Michie-King distribution function, Amorisco & Evans (2012) find 222Although this mass estimator uses the value of the radial profile at , we used the averaged value due to the limited number of stellar tracers of several galaxies in the dataset.. Campbell et al. (2017) used the APOSTLE simulations (Sawala et al., 2016) to derive a mass estimator for larger radii .
For a similar radius, Errani et al. (2018) find that minimises the errors introduced by the shape of the tidally stripped dark matter halo (cusp or core) in which lies a Plummer sphere stellar component.
Some mass estimators have been tested on cosmological simulations. Apart from their own estimator, Campbell et al. (2017) tested Walker’s and Wolf’s mass estimators on fully cosmological hydrodynamical simulations from the APOSTLE project. González-Samaniego et al. (2017) tested Walker’s, Wolf’s, and Campbell’s mass estimators on 10 dwarf galaxy cosmological hydrodynamical simulations from the FIRE-2 project (Hopkins et al., 2018b). The two studies obtained different results for the biases and specially for the scatters of the three mass estimators.